OpenStudy (anonymous):

THE DECIMALS ARE ACTUALLY NEGATIVE EXPONENTS HELP PLEASE??? The weight of a newborn baby chicken weighs 3.2 pounds. If an adult chicken can weigh up to 3.4 times more than a newborn chicken, how much does an adult chicken weigh? Leave your answer in the form ab.

OpenStudy (anonymous):

3.2 * 3.4

OpenStudy (anonymous):

10.88

OpenStudy (anonymous):

multiply the baby weight by the multiplier

OpenStudy (anonymous):

how do I leave it in form ab?

OpenStudy (anonymous):

3.2 * 3.4 is ab

OpenStudy (anonymous):

oh okay, thank you!

OpenStudy (anonymous):

Can you help me with another @peachpi I cant ask another question without verifying me email and I dont have time

OpenStudy (anonymous):

keep going i'm watching

OpenStudy (anonymous):

The world population is increasing about 140 people per minute! At this rate, how much will the population increase in a decade? Put your answer in Scientific Notation. SHOW ALL OF YOUR WORK, THIS INCLUDES USING NUMBERS AND WORDS! Put your answer in Scientific Notation

OpenStudy (anonymous):

Thank you for your help

OpenStudy (anonymous):

um... how many years in a decade?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so... \[3.4\times10\times365\times60~~~ is~the~answer~i~think~\]

OpenStudy (anonymous):

Thank you! This is helping me a bunch

OpenStudy (anonymous):

Can you show me how you got the numbers? I need to show all of my work

OpenStudy (anonymous):

and then try to make it into sciencitific notation so it makes #.##*10^#

OpenStudy (anonymous):

@Leong

OpenStudy (anonymous):

right idea, but you missed a conversion. and I think the 3.4 is for the previous problem. you basically need to find the amount of minutes in a decade and then multiply by 140. 60 minutes in an hour 24 hours in a day 365 day in a year 10 years in a decade

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

and i forgot the 24

OpenStudy (anonymous):

okay, so first I do 24*365=8760 then I 8760*10=87600 Then 87600*140=12,264,000?

OpenStudy (anonymous):

you forgot the 60 minutes

OpenStudy (anonymous):

so I multiply the 60 minutes too? or just mention it?

OpenStudy (anonymous):

no you have to multiply it, because then you have to take the result into scientific notation

OpenStudy (anonymous):

Okay, so all I have to do is 60*24=1440 1440*365=525,600 525,600*10=5,256,000

OpenStudy (anonymous):

um no not like that....

OpenStudy (anonymous):

I wish I was better at math this is really confusing

OpenStudy (anonymous):

yes that's right. there are 5,256,000 minutes in a decade. Now multiply by 140 to find the population increase over the decade

OpenStudy (anonymous):

okay, so 5,256,000*140=735,840,000

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the final answer is 735,840,000? Thank you!

OpenStudy (anonymous):

It says put the answer in scientific notation. Do you know how to do that?

OpenStudy (anonymous):

no, not really :/

OpenStudy (anonymous):

so basicially... 321000000 turns into 3.21*10^8

OpenStudy (anonymous):

Im still confused

OpenStudy (anonymous):

in scientific notation, the decimal needs to be behind the first number that's not 0. If the decimal point moves left, the exponent on 10 is positive. If it moves right the exponent on 10 is negative. The number of the exponent is the number of spaces the decimal moved